OpenStudy (aaronandyson):

three positive number are in the ratio 1/2:1/3:1/4 .Find the numbers if the sum of their squares is 244 @perl

OpenStudy (aaronandyson):

Ratios are" \[\frac{ 1 }{ 2 }:\frac{ 1 }{ 3 }:\frac{ 1 }{ 4 }\]

OpenStudy (aaronandyson):

\[244\] Is the sum of THEIR squares.

OpenStudy (aaronandyson):

@perl

OpenStudy (mathmate):

Hints: 1. sum of squares form an integer, there is a good chance the numbers are integers. 2. numbers are in ratios 1/2:1/3:1/4. The LCM of 2,3,4 is 12. 3. So the numbers are in the ratio 12/2:12/3:12/4=6:4:3. 4. sum of squares of 6,4,3=36+16+9=61 = 244/4. 5. Will let you figure out the rest of the problem.

OpenStudy (aaronandyson):

???

OpenStudy (aaronandyson):

@mathmate

OpenStudy (mathmate):

In math, we do a lot of trial and error, and the deductions made when we get the right (or wrong) answer is the thinking process required in learning math. So I strongly suggest you do some number work to get the answer, or post your tries so we can take it from there.

OpenStudy (perl):

See if you can eliminate y and z, so you have one equation in one variable

OpenStudy (aaronandyson):

???

OpenStudy (perl):

\(\color{blue}{\text{Originally Posted by}}\) @AaronAndyson ??? \(\color{blue}{\text{End of Quote}}\) Can you be more specific

OpenStudy (aaronandyson):

I did not understand what you me to do...

OpenStudy (aaronandyson):

What next @perl?

OpenStudy (aaronandyson):

@welshfella

OpenStudy (perl):

x^2 + y^2 + z^2 = 244 x: y : z = 1/2 : 1/3 : 1/4 multiply ratio through by LCM of 2,3,4 which is 12 1/2 : 1/3 : 1/4 ⟺ 12*1/2 : 12* 1/3 : 12 * 1/4 = 6 : 4: 3 So far we have x^2 + y^2 + z^2 = 244 x: y : z = 6 : 4: 3 Now lets look at the ratio more closely The ratio gives us three equation: x/y = 6/4 -> 4x = 6y x/z = 6/3 -> 3x = 6z y/z = 4/3 -> 3y = 4z

OpenStudy (perl):

Our goal is to produce one equation in one unknown. We have derived the following equations: x^2 + y^2 + z^2 = 244 4x = 6y 3x = 6z 3y = 4z y = 4/3 z x = 6/3 z = 2z so lets substitute these (2z)^2 + (4/3)^2 + z^2 = 244 now we have one equation in one unknown

OpenStudy (aaronandyson):

???

OpenStudy (aaronandyson):

i DIDN'T UNDERSTAND :(

OpenStudy (perl):

what level of math are you, i hope this isn't too advanced

OpenStudy (aaronandyson):

Sophomore in high school Doing this chapter after like MONTHS

OpenStudy (perl):

ok can you point to the line which does not make sense,

OpenStudy (aaronandyson):

The equation you formed...

OpenStudy (perl):

\(\color{blue}{\text{Originally Posted by}}\) @perl x^2 + y^2 + z^2 = 244 x: y : z = 1/2 : 1/3 : 1/4 multiply ratio through by LCM of 2,3,4 which is 12 1/2 : 1/3 : 1/4 ⟺ 12*1/2 : 12* 1/3 : 12 * 1/4 = 6 : 4: 3 So far we have x^2 + y^2 + z^2 = 244 x: y : z = 6 : 4: 3 Now lets look at the ratio more closely The ratio gives us three equation: x/y = 6/4 -> 4x = 6y x/z = 6/3 -> 3x = 6z y/z = 4/3 -> 3y = 4z \(\color{blue}{\text{End of Quote}}\) Does this make sense so far?

OpenStudy (aaronandyson):

Yes.